# Xi+: Uncertainty Supervision for Robust Speaker Embedding

### v1

@@ 1 Introduction
Automatic speaker recognition aims to determine a speaker’s identity solely based on their voice using machine learning algorithms [1]. For each speech utterance, a front-end module extracts a fixed-dimensional speaker embedding that encapsulates speaker-discriminative information while suppressing irrelevant variability such as phonetic content and background noise [2, 3, 4, 5]. Speaker recognition is then performed by computing the similarity between the embeddings of the enrollment and test utterances.
Unlike text or images, speech signals are continuous-valued and variable in length. Moreover, speech conveys not only speaker identity but also a wide range of variations that can degrade the quality of the extracted speaker embeddings. These include linguistic factors (e.g., language), acoustic conditions (e.g., recording channel), and speaker-intrinsic traits (e.g., emotion) [6, 7, 8]. As a result, different temporal frames should contribute unequally to the final fixed-dimensional speaker embedding.
@@ 2.1 Recap of Xi-vector embedding with Uncertainty
Fig. 1: The architecture of xi+, an extended version of the xi-vector model [14]. The modules shown in gray correspond to components of the original xi-vector system, while the modules highlighted in other colors represent our proposed extensions. In the diagram, $\mathbf{z}_{p}$ denotes the prior mean vector, and $\mathbf{L}_{p}$ denotes the prior precision matrix. $\otimes$ and $\ominus$ denotes element-wise multiplication and concatenation, respectively. BNV and FC3V constitute the variance-processing branch, which is structurally parallel to the mean branch. Specifically, they share parameters with BN2 and FC3 in the mean branch, ensuring consistency across the two branches. As shown in distribution space, all frame-level Gaussian distributions are integrated to form an utterance-level Gaussian distribution, which serves as a compact representation of the speaker.
@@
In addition, xi-vector assumes the prior mean and prior covariance of $\mathbf{h}$ as $\mathbf{z}_{p}$ and $\mathbf{L}_{p}^{-1}$ [^1] The prior mean $\mathbf{z}_{p}$ and the prior covariance $\mathbf{L}_{p}^{-1}$ are initialized as 0 and $\mathbf{I}$, respectively, implying that the prior distribution of the latent variable $\mathbf{h}$ is assumed to follow a standard normal distribution.. Mathematically, the posterior distribution of $\mathbf{h}$ is estimated as follows:
@@ 2.2.1 Temporal Modeling
The original xi-vector architecture predicts the precision matrix using fully connected layers, which only capture relationships along the channel dimension. We hypothesize that temporal dependencies are also informative for uncertainty modeling. To this end, we introduce a Transformer encoder containing a single layer with $N$ attention heads within the pooling layer  [^2] NN is set to 8 in this experiment.. This newly added module serves two primary purposes: enhancing the non-linear modeling capacity of the network, and enabling the incorporation of temporal contextual information across frames.
@@ 2.2.2 Stochastic Variance Loss
The key idea of $\mathcal{L}_{\text{SVL}}$ is to establish a pseudo ground truth for uncertainty. Specifically, we regard the distance between a speaker embedding $\boldsymbol{\phi}_{fc}$ and the centroid of embeddings from the same speaker as a pseudo-label. Since each variance estimate is supervised only by a single embedding–centroid pair, rather than the full distribution, we term this approach ‘stochastic variance’ learning.
@@
$\mathbf{c}^{y_{j}}=\frac{\sum_{j=1}^{J}\boldsymbol{\phi}_{fc}^{j}}{J},$ (5)
@@
$\mathcal{L}_{\text{SVL}}=\frac{1}{B}\sum_{b=1}^{B}\|\alpha\sqrt{{(\mathbf{L}_{fc}^{b}})^{-1}}-|\boldsymbol{\phi}_{fc}^{b}-\mathbf{c}^{y_{b}}|\|^{2}_{2}.$ (6)
where $B$ denotes the number of batch size in training. $(\mathbf{L}_{fc}^{b})^{-1}$ and $\mathbf{c}^{y_{b}}$ denote the covariance matrix for utterance $b$ and the centroid of speaker $y_{b}$, respectively. $\|\cdot\|_{2}$ denotes the Euclidean norm. In addition, we introduce a learnable scaling parameter $\alpha$, referred to as the uncertainty scaling factor, to address the potential scale mismatch between the predicted standard deviation $\sqrt{(\mathbf{L}_{fc}^{b})^{-1}}$ and the observed embedding deviation $|\boldsymbol{\phi}_{fc}^{b}-\mathbf{c}^{y_{b}}|$.
@@
$\boldsymbol{\phi}_{bn}=\frac{\boldsymbol{\phi}-\boldsymbol{\mu}_{bn}}{\sqrt{\boldsymbol{\sigma}_{bn}+\boldsymbol{\epsilon}}}\otimes\boldsymbol{\gamma}+\boldsymbol{\beta},$ (7)
where $\boldsymbol{\mu}_{bn}$ and $\boldsymbol{\sigma}_{bn}$ denote the batch-wise mean and variance, respectively. $\boldsymbol{\epsilon}$ is a small constant for numerical stability, and $\boldsymbol{\gamma}$ and $\boldsymbol{\beta}$ are learnable affine parameters. The symbol $\otimes$ denotes element-wise multiplication. Subsequently, the normalized vector is passed through a fully connected layer:
$\boldsymbol{\phi}_{fc}=\boldsymbol{\phi}_{bn}\mathbf{W}^{\top}+\mathbf{b},$ (8)
@@
$\mathbf{L}_{bn}^{-1}=\frac{\mathbf{L}^{-1}\otimes\mathbf{\boldsymbol{\gamma}}^{2}}{\boldsymbol{\sigma}_{bn}+\boldsymbol{\epsilon}}.$ (9)
@@
$\mathbf{L}_{fc}^{-1}=\mathbf{W}\mathbf{L}_{bn}^{-1}\mathbf{W}^{\top}.$ (10)
@@ 2.2.3 Cosine Scoring with Uncertainty
$s_{u}=\frac{<\boldsymbol{\phi}^{s2}_{fc},\boldsymbol{\phi}^{s1}_{fc}>}{\sqrt{\boldsymbol{\phi}{{}^{s2}_{fc}}^{\top}(\mathbf{I}+\rho(\mathbf{L}_{fc}^{s2})^{-1})^{-1}\boldsymbol{\phi}^{s2}_{fc}}\sqrt{\boldsymbol{{\phi}}{{}^{s1}_{fc}}^{\top}(\mathbf{I}+\rho(\mathbf{L}^{s1}_{fc})^{-1})^{-1}\boldsymbol{\phi}^{s1}_{fc}}}.$ (11)
@@ 3.1 Training details
where $\kappa$ is used to control loss weight and is dynamically adjusted based on the current epoch $ep$ to ensure that $\mathcal{L}_{\text{SVL}}$ gradually contributes as training progresses:
@@ 3.2 Dataset
We conduct two sets of experiments. The first set is performed on Voxceleb2, which consists of 5,994 speakers, and is evaluated on Voxceleb1 [22, 23, 24]. The second set is conducted on the NIST SRE 2024 dataset. We follow the evaluation plan outlined in [25] to construct our training, development, and evaluation datasets. Additionally, we apply rVAD [26] to remove non-speech signals. For both sets, we also employ data augmentation techniques, incorporating noise from the MUSAN database [27] and room impulse responses from the RIR database [28].
@@ 4.1 Results on Voxceleb
Table 2: Comparison of different systems employing cosine scoring on on SRE 2024 development and evaluation sets in terms of EER. xi+ adopts the best configuration from Table 1, which is highlighted in gray. ‘LM’ denotes large margin finetune. In this table, a new speaker encoder Golden Gemini [29] is adopted.

### v2

@@ 1 Introduction
Automatic speaker recognition aims to determine a speaker’s identity solely based on their voice using machine learning algorithms [1]. For each speech utterance, a front-end module extracts a fixed-dimensional speaker embedding that encapsulates speaker-discriminative information while suppressing irrelevant variability such as phonetic content and background noise [2, 3, 4, 5, 6]. Speaker recognition is then performed by computing the similarity between the embeddings of the enrollment and test utterances.
Unlike text or images, speech signals are continuous-valued and variable in length. Moreover, speech conveys not only speaker identity but also a wide range of variations that can degrade the quality of the extracted speaker embeddings. These include linguistic factors (e.g., language), acoustic conditions (e.g., recording channel), and speaker-intrinsic traits (e.g., emotion) [7, 8]. As a result, different temporal frames should contribute unequally to the final fixed-dimensional speaker embedding.
@@ 2.1 Recap of Xi-vector embedding with Uncertainty
[…]
@@
In addition, xi-vector assumes the prior mean and prior covariance of $\mathbf{h}$ as $\mathbf{z}_{p}$ and $\mathbf{L}_{p}^{-1}$ [^1] The prior mean $\mathbf{z}_{p}$ and the prior covariance $\mathbf{L}_{p}^{-1}$ are learnable parameters and initialized as 0 and $\mathbf{I}$, respectively, implying that the prior distribution of the latent variable $\mathbf{h}$ is assumed to follow a standard normal distribution.. Mathematically, the posterior distribution of $\mathbf{h}$ is estimated as follows:
@@ 2.2.1 Temporal Modeling
The original xi-vector architecture predicts the precision matrix using fully connected layers, which only capture relationships along the channel dimension. We hypothesize that temporal dependencies are also informative for uncertainty modeling. To this end, we introduce a Transformer encoder containing a single layer with NN attention heads within the pooling layer  [^2] NN is set to 8 in this experiment.. This newly added module serves two primary purposes: enhancing the non-linear modeling capacity of the network, and enabling the incorporation of temporal contextual information across frames.
@@ 2.2.2 Stochastic Variance Loss
The key idea of $\mathcal{L}_{\text{SVL}}$ is to establish a pseudo ground truth for uncertainty. Specifically, we regard the distance between a speaker embedding $\boldsymbol{\phi}_{\text{fc}}$ and the centroid of embeddings from the same speaker as a pseudo-label. Since each variance estimate is supervised only by a single embedding–centroid pair, rather than the full distribution, we term this approach ‘stochastic variance’ learning.
@@
$\mathbf{c}^{y_{j}}=\frac{\sum_{j=1}^{J}\boldsymbol{\phi}_{\text{fc}}^{j}}{J},$ (5)
@@
$\mathcal{L}_{\text{SVL}}=\frac{1}{B}\sum_{b=1}^{B}\|\alpha\sqrt{{(\mathbf{L}_{\text{fc}}^{b}})^{-1}}-|\boldsymbol{\phi}_{\text{fc}}^{b}-\mathbf{c}^{y_{b}}|\|^{2}_{2},$ (6)
where $B$ denotes the number of batch size in training. $(\mathbf{L}_{\text{fc}}^{b})^{-1}$ and $\mathbf{c}^{y_{b}}$ denote the covariance matrix for utterance $b$ and the centroid of speaker $y_{b}$, respectively. $\|\cdot\|_{2}$ denotes the Euclidean norm. In addition, we introduce a learnable scaling parameter $\alpha$, referred to as the uncertainty scaling factor, to address the potential scale mismatch between the predicted standard deviation $\sqrt{(\mathbf{L}_{\text{fc}}^{b})^{-1}}$ and the observed embedding deviation $|\boldsymbol{\phi}_{\text{fc}}^{b}-\mathbf{c}^{y_{b}}|$.
@@
$\boldsymbol{\phi}_{\text{bn}}=\frac{\boldsymbol{\phi}-\boldsymbol{\mu}_{\text{bn}}}{\sqrt{\boldsymbol{\sigma}_{\text{bn}}+\boldsymbol{\epsilon}}}\otimes\boldsymbol{\gamma}+\boldsymbol{\beta},$ (7)
where $\boldsymbol{\mu}_{\text{bn}}$ and $\boldsymbol{\sigma}_{\text{bn}}$ denote the batch-wise mean and variance, respectively. $\boldsymbol{\epsilon}$ is a small constant for numerical stability, and $\boldsymbol{\gamma}$ and $\boldsymbol{\beta}$ are learnable affine parameters. The symbol $\otimes$ denotes element-wise multiplication. Subsequently, the normalized vector is passed through a fully connected layer:
$\boldsymbol{\phi}_{\text{fc}}=\boldsymbol{\phi}_{\text{bn}}\mathbf{W}^{\top}+\mathbf{b},$ (8)
@@
$\mathbf{L}_{\text{bn}}^{-1}=\frac{\mathbf{L}^{-1}\otimes\mathbf{\boldsymbol{\gamma}}^{2}}{\boldsymbol{\sigma}_{\text{bn}}+\boldsymbol{\epsilon}}.$ (9)
@@
$\mathbf{L}_{\text{fc}}^{-1}=\mathbf{W}\mathbf{L}_{\text{bn}}^{-1}\mathbf{W}^{\top}.$ (10)
@@ 2.2.3 Cosine Scoring with Uncertainty
$s_{u}=\frac{<\boldsymbol{\phi}^{s2}_{\text{fc}},\boldsymbol{\phi}^{s1}_{\text{fc}}>}{\sqrt{\boldsymbol{\phi}{{}^{s2}_{\text{fc}}}^{\top}(\mathbf{I}+\rho(\mathbf{L}_{\text{fc}}^{s2})^{-1})^{-1}\boldsymbol{\phi}^{s2}_{\text{fc}}}\sqrt{\boldsymbol{{\phi}}{{}^{s1}_{\text{fc}}}^{\top}(\mathbf{I}+\rho(\mathbf{L}^{s1}_{\text{fc}})^{-1})^{-1}\boldsymbol{\phi}^{s1}_{\text{fc}}}}.$ (11)
@@ 3.1 Training details
where $\kappa$ is used to control loss weight and is dynamically adjusted based on the current epoch epep to ensure that $\mathcal{L}_{\text{SVL}}$ gradually contributes as training progresses:
@@ 3.2 Dataset
We conduct two sets of experiments. The first set is performed on Voxceleb2, which consists of 5,994 speakers, and is evaluated on Voxceleb1 [22]. The second set is conducted on the NIST SRE 2024 dataset. We follow the evaluation plan outlined in [23] to construct our training, development, and evaluation datasets. Additionally, we apply rVAD [24] to remove non-speech signals. For both sets, we also employ data augmentation techniques, incorporating noise from the MUSAN database [25] and room impulse responses from the RIR database [26].
@@ 4.1 Results on Voxceleb
Table 2: Comparison of different systems employing cosine scoring on on SRE 2024 development and evaluation sets in terms of EER. xi+ adopts the best configuration from Table 1, which is highlighted in gray. ‘LM’ denotes large margin fine-tuning [27]. In this table, a new speaker encoder Golden Gemini [28] is adopted for all models.